OpenStudy (anonymous):

Find the amplitude and period of f(t)=-tan 0.4t

OpenStudy (kc_kennylau):

This may help you: http://answers.yahoo.com/question/index?qid=20080819015140AAP4IAx

OpenStudy (campbell_st):

well the tan function doesn't have an amplitude, as it contains vertical asymptotes the period for y=tan(bx) is \[\frac{\pi}{b}\] in your question you have b = 0.4 you should leave your answer in terms of pi... hope this helps